A cylindrical container has a base area of 100 M and is 12 m hight. If the container is one-third filled with water, what is the volume of the water in the container?
Answer by Theo(13342)

volume of a cylinder = pi*r^2*h

r is the radius.
h is the height.

base area of the cylinder = pi*r^2 = 100 square meters.

height of the cylinder = h = 12 meters.

volume of the cylinder = pr^2 * h = 100 * 12 = 1200 cubic meters.

Since the cylinder is 1/3 filled with water, this means that the volume of the water is 1/3 the volume of the cylinder which means that the volume of the water is 400 cubic meters.

Another way of looking at this is that the height of the cylinder is 12 meters and it is 1/3 filled with water which means that the height of the water is 4 meters.

4 * 100 = 400 cubic meters.

This is the same result looking at it a different way.

Your answer is that the volume of the water is 400 cubic meters.